# Designing Library of Skill-Agents for Hardware-Level Reusability

### 04 Mar 2024 v1

@@ 4.3.2 OP-PR (Drawer-open), PR-PR (Drawer adjust), PR-OP (Drawer-close) skills
All the skills need to satisfy the same rule. Generally, the skill parameters (e.g., drawing direction) observed in the demonstration include errors. Thus, it is necessary to train the skills using RL. In order to reduce the training effort, we train the PR-PR skill first, and then add the program to decide the terminal condition.
The state in PR can be represented by setting the direction of the feasible displacement. The skill needs the force information for feedback in the execution. Thus, the state in RL can be designed as follows:
@@
$\mathbf{c}_{t+1}=\frac{\mathbf{c}_{t}+\Delta c}{|\mathbf{c}_{t}+\Delta c|}.$
The PR-PR skill requires to satisfy F-t $\leq$ delta-collision and F-u $\leq$ delta-collision. If the force exerted by the constraint would be reduced, these two conditions tend to be satisfied. Thus, the reward function $r$ can be formulated as follows:
$r=-|f_{s}|$ (4)
The episode in RL training is terminated after the predefined duration. Figure 31 shows the reward curve. The training was terminated after one million steps. The reward curve looks to converge.
Figure 31: Reward curve in PR-PR
Figure 32: Execution of PR-OP
Figure 33: Force profile along x-axis (displacement direction) in PR-OP
Figure 34: Force profiles along y-axis (horizontal direction) and z-axis (vertical direction) in PR-OP
Figure 32 shows the execution of the PR-OP skill by an actual robot. We set the distance of the displacement at each step to 5 [mm]. Figure 33 shows the force along the drawing direction. Closing the drawer generates a large reaction force, which allows us to determine the termination of the PR-OP skill. This drawer has a locking mechanism that prevents it from opening accidentally on its own. It can be seen that a large force is generated at time 21 before the lock, then the force decreases, and finally a large force is generated when the drawer is completely closed. Figure 34 shows the changes in the force values in the horizontal and vertical directions, orthogonal to the direction of drawing. Although there are some situations where unwanted force is generated during the execution, the closing direction is modified to suppress the generation of force in the PR-OP skill. However, instability is confirmed because of the deviation from the physical condition of PR-PR around the time it approaches the locking mechanism (See force along the z-axis), but PR-OP is achieved without any troubles because of the short time from there to the closure.
4.3.3 OR-RV (Door-open), RV-OR (Door-close), RV-RV (Door-adjust) skills
These skills can be described by the following control rule without the termination condition.
@@ 4.3.3 OR-RV (Door-open), RV-OR (Door-close), RV-RV (Door-adjust) skills
All the skills need to satisfy this rule. Generally, the skill parameters (e.g., the configuration of the rotation axis) observed in the demonstration include errors. These skills also need to be trained using RL. As described above, the infinitesimal displacement at each moment can also be interpreted as an infinitesimal translation tangential to the rotation. The whole trajectory can be regarded as the pieces of the infinitesimal translation, where the translation direction is gradually changed. We can regard that the training of RV-RV is the same as that of PR-PR, since the target displacement at both skills needs to be modified following the constraint (e.g., force feedback) of the (infinitesimal) translation; the control rule is the same.
@@
The action is also the same as that in PR-PR as $\Delta\mathbf{c}\equiv(\Delta c_{x},\Delta c_{y},\Delta c_{z})$. The modification is also the same as
$\mathbf{c}_{t+1}=\frac{\mathbf{c}_{t}+\Delta c}{|\mathbf{c}_{t}+\Delta c|}.$
The difference between RV-RV and PR-PR is that the RV-RV involves the orientation changes with respect to the modification of the displacement. Thus, we add the orientation changes to the PR-PR skill; if the displacement is changed by $\theta$, the orientation of the hand, as well as the next motion direction, are modified by $\theta$ around the center of the grasping points. See Figure 35.
Figure 35: RV-RV transition. Q1 and Q2 are the contact points at each instant, and S is the axis of rotation. The motion direction should be corrected at each regular interval.
Figure 36: Execution of OR-RV
@@
Figure 36 shows the execution of the OR-RV skill by an actual robot. In this case, we also set the distance of the displacement at each step to 5 [mm]. Figure 37 shows the trajectory of the center of the grasping points and the estimated opening directions. As can be seen, the robot tried to follows the curvilinear trajectory to open the door and the estimated opening direction is modified to follow the tangential direction of the trajectory.
4.3.4 PC1-PC1 (Wipe) skill
@@
From the control rule, the force along the normal direction, F-u, should be less than delta-collision and more than delta-zero. Then, the PC1-PC1 skill is realized by controlling F-u to be an appropriate value $f_{c}$ in between. Specifically, if the value of the force sensor when contacting each other is $\mathbf{f}_{0}$, the target force $\mathbf{f}_{d}$ is calculated by the following equation:
@@
  • target displacement direction: $\Delta\mathbf{d}$
  • Normalized difference between current and target forces: $f_{s}\equiv\Delta\mathbf{f}_{t}/|\Delta\mathbf{f}_{t}|$,
where $\Delta\mathbf{f}_{t}\equiv\mathbf{f}_{t}-\mathbf{f}_{0}$. Unlike the PR-PR and RV-RV cases, which merely reduce the norm of the force, the PC1-PC1 skill need to consider the magnitude of the force in order to control F-u to approach to $f_{c}$. Therefore, a coarse discretization of the magnitude of $\Delta\mathbf{f}_{t}$ is also included as a state. In other words, we add to the state the following element:
  • The information on the magnitude of the force: $\lfloor|\Delta\mathbf{f}_{t}|/f_{step}\rfloor$.
@@
$r=\left\{\begin{array}[]{ll}-f_{max}&(f_{s}>f_{max})\\ -f_{max}&(\hbox{the contact is detached})\\ f_{max}/2-f_{s}&(\hbox{otherwise})\end{array}\right.$ (6)
@@
Figure 38: Reward curve in training PC1-PC1
Figure 38 shows the reward curve. We finished training after two million steps. The reward looks to be converged.
Figure 39: Execution of PC1-PC1
Figure 40: Force profile along z-axis in PC1-PC1
@@
Figure 39 shows the execution of the trained PC1-PC1 skill by an actual robot. The displacement for each step is 5 [mm] in the target translational direction plus the modification output by the trained RL agent. It can be seen that the motion is modified to achieve the target of 10 [N]. As shown in Figure 41, the PC1-PC1 skill, i.e., wiping, was actually able to erase the drawing on the white board. Theoretically, as $f_{step}$ is increased, the degree of feedback to the force value can be reduced. Further improvement of control may be possible by adjusting $f_{step}$.
@@ 5 Working system
This section describes how the implemented skills function in a reusable manner within the end-to-end Learning-from-observation system.
@@ 5.1 Observation station
Figure 42 shows the observation station. For observation during the demonstration, we utilized an RGB-D camera, specifically Azure Kinect by Microsoft. To ensure alignment in the orientation between the robot coordinates and the demonstration coordinates, we employed an AR marker. This alignment enables the robot to replicate the demonstrated task sequence, achieving the same displacement while incorporating collision avoidance measures. Note that the locations of the objects do not have to be exactly the same, although they are assumed to be approximately the same during the demonstration and runtime. Any small differences are accommodated by the respective skills.
@@
From the explanation and the demonstration image sequences, the tasks and their skill parameters can be estimated (Wake et al. (2021)). Initially, the type of each task is estimated. Subsequently, the skill parameters for each task are determined by observing the task sequence again. Typically, the skill parameters encompass the hand position and Labanotation pose at the beginning and end of the task. Additionally, in grasping, the approach direction are also included as the skill parameters.
@@ 5.2 Hand motion to body motion under hardware-level reusability
Using IK, we transform hand motions outputted by the skill into body motions. A typical IK solver, such as Beeson and Ames (2015), minimizes the difference between the desired hand pose and the pose of the hand obtained through forward kinematics under a certain joint configuration. Generally, there are multiple solutions that satisfy the target hand pose and IK sometimes outputs unexpected body pose. As the degrees of freedom of a robot increase, the likelihood of unexpected poses occurring becomes more pronounced. This issue is addressed by using a Labanotation-based IK solver with the body-role division algorithm (Sasabuchi et al. (2021)).
@@ 5.3 Robot testbed
We used three testbed robots shown in Figure 44. Both robots run on ROS (Quigley (2009)). The first (refered to as Nextage-Shadow) was Nextage, Kawada Robotics [^4] https://www.kawadarobot.co.jp/en/nextage/. It has two arms and each arm has six DOF. It also has one DOF in the waist (rotation around the vertical axis). In this paper, we used only the right arm, not the left arm or waist, to perform tasks. The right arm was equipped with 6-axis force/torque sensor, FFS Series, Leptrino [^5] https://www.leptrino.co.jp/product/6axis-force-sensor (Japanese) and Shadow Dexterous Hand Lite, Shadow Robotics [^6] https://www.shadowrobot.com/dexterous-hand-series/, as a robot hand. Nextage was equipped with a stereo camera to observe an environment in a 3-dimensional space.
@@ 6 Demonstration
To demonstrate hardware-level reusability in the learning-from-observation framework, we applied the four same task sequences, Place-on-plate demo, shelf-sequence demo, throw-away demo, and open-fridge demo to the three different robots, Nextage-Shadow, Fetch-and Fetch-Parallel. The place-on-plate demo consists of grasping the box, picking up the box from a table, bringing the box, placing the box on a plate, and releasing the box (See Figure 43). The shelf-sequence demo consists of grasping the cup, picking up the cup, three repetitions of bringing the cup, and releasing the cup (See Figure 46). The throw-away demo consists of picking up the cup, bringing the cup, and releasing the box (See Figure 47). The open-fridge demo consists of grasping the handle and opening the fridge (see Figure 48). All the execution videos can be seen from https://j-taka.github.io/research/hardware_level_reusability.html.
@@ 6.3 Throw-away demo
Figure 48: Open-fridge demo. The first row: Human demonstration. The second row: Execution by Nextage-Shadow. The third row: Execution by Fetch-Shadow. The fourth row: Execution by Fetch-Parallel
@@ 6.5.1 How reusable?
Since the hand is the same in Nextage-Shadow and Fetch-Shadow, the implemented skills (both manipulation and grasping skills) are reusable by changing the IK solver for Nextage or Fetch. Since the hand is different in Fetch-Shadow and Fetch-Parallel, it is necessary to change the grasping skills. Fortunately, the parallel gripper has only one DOF, just opening and closing, and thus, the grasping is achieved by positioning the hand and closing. That is not so difficult to implement. Furthermore, the manipulation skills can be used as is; the finger joints are fixed in the execution of these skills and the execution is completed by body motion.
Through the experiments, we found the following two things. First, we found that each hardware has a different range of acceptable skill parameters. For example, in the place-on-plate demo, the approach direction of the parallel gripper is changed to the upward to increase the success rate of the task sequence. Conversely, if the acceptable range is known, the success rate can be increased by adjusting the parameters without changing the nature of the target task sequence. Second, we found that it may be better not to use human grasping strategies as they are in the parallel gripper, because the parallel gripper differs significantly from a human hand in the shape and DOF. OR-RV (PTG51) can be achieved under satisfying the assumption where a hand and an object are integrated into one unit. Though Fetch-Parallel was able to open the fridge, the excessive feedback may be applied due to the deviation from the condition at the beginning of the opening. Although Shadow Dexterous Hand Lite can be hooked while wrapping the handle, the parallel gripper touch the handle at small number of contact points.
@@ 6.5.2 Comparison to related work
On the other hand, the target user of the proposed system is an ordinary person, who does not have a programming skill. For such a person, it is desirable to enable to make a program for place-to-plate demo using the indications, such as grasping a box and pick it up. That requires the robot system to remove the effort of programming such as use of the vision and mathematics to decide the pose of the gripper. The proposed system provides those using the hint from the human demonstration. Using more abstracted instructions can remove awareness of the hardware. For example, issues related to approach direction in grasping are solved by the system.
@@ 7 Summary and discussion
This paper focused on the second step of Learning-from-observation (LfO), emphasizing the portion where the task models, obtained from demonstrations, were executed on the various robots. In line with the spirit of LfO, an effort is made to create a system that is as independent on the robot’s hardware as possible. Attention was directed towards hand motions to design skills that move the hand to satisfy conditions derived from the task requirements and physical constraints from the environment. By taking hand motions as a reference, differences in hardware among robots are absorbed into inverse kinematics (IK), allowing the easy substitution of an appropriate IK when hardware changes, ensuring the same skills can be used.
In environments with physical constraints, adjustments to hand trajectories are often necessary; such a concept is known as compliant manipulation. Recently, reinforcement learning (RL) has been applied to address household tasks, including compliant manipulation. However, traditional RL methods have primarily focused on designing policies for specific tasks, limiting their applicability and requiring individual learning for each new task. We proposed policies that consider constraints applicable to various tasks by grouping multiple tasks based on the types of physical constraints. The type of physical constraints are determined by the characteristics of the imposed force directions. Consequently, a general policy is learned based on these characteristics.
End-to-end generation of robot programs using large language models (LLM) has been proposed in general, such as Vemprala et al. (2023); Wake et al. (2023); Yu et al. (2023). However, many robot tasks require compliant manipulation. In the case of compliant manipulation, fine-tuning the hand trajectory using force feedback becomes necessary. Therefore, while systems based on LLMs can provide an overview of program design, generating programs with fine-tuning capabilities using local feedback is challenging. To address this issue, we propose first preparing a standard set of machine-independent executable skills using reinforcement learning or similar methods. This standard set is then provided as a prompt to the LLMs, enabling the design of a robot system that can actually execute. We consider this approach holds promise for resolving the challenge of incorporating local feedback in robot program generation using LLMs.
In this paper, we utilized the Shadow Hands on two different robots. As a result, we were able to share the same grasp skill library. In the case of the parallel gripper of Fecth, we resolved the difference in the hardware by only changing the grasp skill library. Standardizing grasp skills across different robotic hands remains a challenge for future research.

### 20 Mar 2024 v2

@@ 4.3.2 OP-PR (Drawer-open), PR-PR (Drawer-adjust), PR-OP (Drawer-close) skill agents
All the skills need to satisfy the same rule. Generally, the skill parameters (e.g., drawing direction) observed in the demonstration include errors. Thus, it is necessary to train these skill agents using RL. In order to reduce the training effort, we train the PR-PR skill agent first, and then add the program to decide the terminal condition.
The state in PR can be represented by setting the direction of the feasible displacement. The skill agnet needs the force information for feedback in the execution. Thus, the state in RL can be designed as follows:
@@
$\mathbf{c}_{t+1}=\frac{\mathbf{c}_{t}+\Delta\mathbf{c}}{|\mathbf{c}_{t}+\Delta\mathbf{c}|}.$
The PR-PR skill agent requires to satisfy F-t $\leq$ delta-collision and F-u $\leq$ delta-collision. If the force exerted by the constraint would be reduced, these two conditions tend to be satisfied. Thus, the reward function $r$ can be formulated as follows:
$r=-|\mathbf{f}_{t}|.$ (4)
The episode in RL training is terminated after the predefined duration. Figure 31 shows the reward curve. The training was terminated after one million steps. The reward curve looks converging.
Figure 31: Reward curve in PR-PR skill agent
Figure 32: Execution of PR-OP (Drawer-close) skill agent
Figure 33: Force profile along x-axis (drawing direction) in PR-OP skill
Figure 34: Force profiles along y-axis (horizontal direction) and z-axis (vertical direction) in PR-OP skill
Figure 32 shows the execution of the PR-OP skill agent by an actual robot. We set the distance of the displacement at each step to 5 [mm]. Figure 33 shows the force along the drawing direction. Closing the drawer generates a large drag force, which allows us to determine the termination of the PR-OP skill. This drawer has a locking mechanism that prevents it from opening accidentally on its own. It can be seen that a large force is generated at time 21 before the lock, then the force decreases, and finally a large force is generated when the drawer is completely closed. Figure 34 shows the changes in the force values in the horizontal and vertical directions, orthogonal to the drawing direction. Although there are some situations where unwanted force is generated during the execution, the drawing direction is modified to suppress the generation of force by the PR-OP skill agent. However, instability is confirmed because of the deviation from the physical condition of PR-PR around the time it approaches the locking mechanism (See force along the z-axis), but the PR-OP skill agent achieved the task without any troubles because of the short time from there to the closure.
4.3.3 OR-RV (Door-open), RV-OR (Door-close), RV-RV (Door-adjust) skill agents
These skill agents can be described by the following control rule without the termination condition.
@@ 4.3.3 OR-RV (Door-open), RV-OR (Door-close), RV-RV (Door-adjust) skill agents
All the skills need to satisfy this rule. Generally, the skill parameters (e.g., the configuration of the rotation axis) observed in the demonstration include errors. These skill agents also need to be trained using RL. As described above, the infinitesimal displacement at each moment can also be interpreted as an infinitesimal translation tangential to the rotation. The whole trajectory can be regarded as the pieces of the infinitesimal translation, where the translation direction is gradually changed. We can regard that the training of RV-RV is the same as that of PR-PR, since the target displacement at both skills needs to be modified following the constraint (e.g., force feedback) of the (infinitesimal) translation; the control rule is the same.
@@
The action is also the same as that in the PR-PR skill agent as $\Delta\mathbf{c}\equiv(\Delta c_{x},\Delta c_{y},\Delta c_{z})$. The modification is also the same as
$\mathbf{c}_{t+1}=\frac{\mathbf{c}_{t}+\Delta\mathbf{c}}{|\mathbf{c}_{t}+\Delta\mathbf{c}|}.$
The difference between the RV-RV and PR-PR skills is that the RV-RV skill involves the orientation changes with respect to the modification of the displacement. Thus, we add the orientation changes to the PR-PR skill agent; if the displacement is changed by $\theta$, the orientation of the hand, as well as the next motion direction, are modified by $\theta$ around the center of the grasping points. See Figure 35.
Figure 35: RV-RV transition. Q1 and Q2 are the contact points at each instant, and S is the axis of rotation. The motion direction and orientation should be corrected at each regular interval.
Figure 36: Execution of OR-RV (Door-opening) skill agent
@@
Figure 36 shows the execution of the OR-RV skill agent by an actual robot. In this case, we also set the distance of the displacement at each step to 5 [mm]. Figure 37 shows the trajectory of the center of the grasping points and the estimated opening directions. As can be seen, the robot tried to follows the curvilinear trajectory to open the door and the estimated opening direction is modified to follow the tangential direction of the trajectory.
4.3.4 PC1-PC1 (Wipe) skill agent
@@
From the control rule, the force along the normal direction, F-u, should be less than delta-collision and more than delta-zero. Then, the PC1-PC1 skill is realized by controlling F-u to be an appropriate value $f_{c}$ in between. Specifically, if the value of the force sensor just when contacting each other is $\mathbf{f}_{0}$, the target force $\mathbf{f}_{d}$ is calculated by the following equation:
@@
  • target displacement direction: $\Delta\mathbf{d}$,
  • Normalized difference between current and target forces: $\mathbf{f}_{n}\equiv\Delta\mathbf{f}_{t}/|\Delta\mathbf{f}_{t}|$,
where $\Delta\mathbf{f}_{t}\equiv\mathbf{f}_{t}-\mathbf{f}_{0}$. Unlike the cases of the PR-PR and RV-RV skill agents, which merely reduce the norm of the force, the PC1-PC1 skill agents need to consider the magnitude of the force in order to control F-u to approach to $f_{c}$. Therefore, a coarse discretization of the magnitude of $\Delta\mathbf{f}_{t}$ is also included as a state. In other words, we add to the state the following element:
  • The information on the magnitude of the force: $f_{desk}\equiv\lfloor|\Delta\mathbf{f}_{t}|/f_{step}\rfloor$.
@@
$r=\left\{\begin{array}[]{ll}-f_{max}&(f_{desk}>f_{max})\\ -f_{max}&(\hbox{the contact is detached})\\ f_{max}/2-f_{desk}&(\hbox{otherwise})\end{array}\right..$ (6)
@@
Figure 38: Reward curve in PC1-PC1 skill agent
Figure 38 shows the reward curve. We finished training after two million steps. The reward looks converging.
Figure 39: Execution of PC1-PC1 (Wipe) skill agents
Figure 40: Force profile along z-axis in PC1-PC1 skill
@@
Figure 39 shows the execution of the PC1-PC1 skill agent by an actual robot. The displacement for each step is 5 [mm] in the target translational direction plus the modification outputted by the skill agent. It can be seen that the motion is modified to achieve the target of 10 [N]. As shown in Figure 41, the PC1-PC1 (Wipe) skill agent was actually able to erase the drawing on the white board. Theoretically, as $f_{step}$ is increased, the degree of feedback to the force value can be reduced. Further improvement of control may be possible by adjusting $f_{step}$.
@@ 5 Working system
This section describes how the implemented skill agents function in a reusable manner within the end-to-end Learning-from-observation system.
@@ 5.1 Observation station
Figure 42 shows the observation station. For observation during the demonstration, we utilized an RGB-D camera, specifically Azure Kinect by Microsoft. To ensure alignment in the orientation between the robot coordinates and the demonstration coordinates, we employed an AR marker. This alignment enables the robot to replicate the demonstrated task sequence, achieving the same displacement while incorporating collision avoidance measures. Note that the locations of the objects do not have to be exactly the same, although they are assumed to be approximately the same during the demonstration and runtime. Any small differences are accommodated by the respective skill agents.
@@
From the explanation and the demonstration image sequences, the tasks and their skill parameters can be estimated (Wake et al. (2021)). Initially, the type of each task is estimated. Subsequently, the skill parameters for each task are determined by analyzing the task sequence again using the daemon functions.
@@ 5.2 Hand motion to body motion under hardware-level reusability
Using IK, we transform hand motions outputted by the skill agent into body motions. A typical IK solver, such as Beeson and Ames (2015), minimizes the difference between the desired hand pose and the pose of the hand obtained through forward kinematics under a certain joint configuration. Generally, there are multiple solutions that satisfy the target hand pose and IK sometimes outputs unexpected body pose. As the degrees of freedom of a robot increase, the likelihood of unexpected poses occurring becomes larger. This issue is addressed by using a Labanotation-based IK solver with the body-role division algorithm (Sasabuchi et al. (2021)).
@@ 5.3 Robot testbed
We used three testbed robots shown in Figure 44. Both robots run on ROS (Quigley (2009)). The first (refered to as Nextage-Shadow) was Nextage, Kawada Robotics [^4] https://www.kawadarobot.co.jp/en/nextage/. It has two arms and each arm has six DOF. It also has one DOF in the waist (rotation around the vertical axis). In this paper, we used only the right arm, neither the left arm nor waist, to perform tasks. The right arm was equipped with 6-axis force/torque sensor, FFS Series, Leptrino [^5] https://www.leptrino.co.jp/product/6axis-force-sensor (Japanese) and Shadow Dexterous Hand Lite, Shadow Robotics [^6] https://www.shadowrobot.com/dexterous-hand-series/, as a robot hand. Nextage was equipped with a stereo camera to observe an environment in a 3-dimensional space.
@@ 6 Demonstration
To demonstrate hardware-level reusability in the learning-from-observation framework, we applied the four task sequences, place-on-plate demo, shelf-sequence demo, throw-away demo, and open-fridge demo to the three different robots, Nextage-Shadow, Fetch-Shadow, and Fetch-Parallel. The place-on-plate demo consists of grasping the box, picking up the box from a table, bringing the box, placing the box on a plate, and releasing the box (See Figure 43). The shelf-sequence demo consists of grasping the cup, picking up the cup, three repetitions of bringing the cup, and releasing the cup (See Figure 46). The throw-away demo consists of picking up the red can, bringing the can, and releasing the can (See Figure 47). The open-fridge demo consists of grasping the handle and opening the fridge (see Figure 48). All the execution videos can be seen from https://j-taka.github.io/research/hardware_level_reusability.html.
@@ 6.3 Throw-away demo
Figure 48: Open-fridge demo. The first row: human demonstration. The second row: execution by Nextage-Shadow. The third row: execution by Fetch-Shadow. The fourth row: execution by Fetch-Parallel.
@@ 6.5.1 How reusable?
Since the hand is the same in Nextage-Shadow and Fetch-Shadow, the implemented skill agents (both manipulation and grasping skill agents) are reusable by changing the IK solver for Nextage or Fetch. Since the hand is different in Fetch-Shadow and Fetch-Parallel, it is necessary to change the grasping skill agents. Fortunately, the parallel gripper has only one DOF, just opening and closing, and thus, the grasping is achieved by positioning the hand and closing. That is not so difficult to implement. Furthermore, the manipulation skill agents can be used as is; the finger joints are fixed in the execution of these skills and the execution is completed by body motion.
Through the experiments, we found the following two things. First, we found that each hardware has a different range of acceptable skill parameters. For example, in the place-on-plate demo, the approach direction of the parallel gripper is changed to the upward to increase the success rate of the task sequence. Conversely, if the acceptable range is known, the success rate can be increased by adjusting the parameters without changing the nature of the target task sequence. Second, we found that it may be better not to use human grasping strategies as they are in the parallel gripper, because the parallel gripper differs significantly from a human hand in the shape and DOF. The OR-RV (PTG51) task can be achieved under satisfying the assumption where a hand and an object are integrated into one unit. Though Fetch-Parallel was able to open the fridge, the excessive feedback may be applied due to the deviation from the assumption at the beginning of the opening. Although Shadow Dexterous Hand Lite can be hooked while wrapping the handle, the parallel gripper touch the handle at small number of contact points.
@@ 6.5.2 Comparison to related work
On the other hand, the target user of the proposed system is an ordinary person, who does not have a programming skill. For such a person, it is desirable to enable to make a program for the place-to-plate demo using the instruction, such as grasping a box and pick it up. That requires the robot system to remove the effort of programming such as use of the vision and mathematics to decide the pose of the gripper. The proposed system provides those using the hint from the human demonstration. Using more abstracted instructions can remove awareness of the hardware. For example, issues related to approach direction in grasping are solved by the system.
@@ 7 Summary and discussion
This paper focused on the second step of Learning-from-observation (LfO), emphasizing the portion where the task models, obtained from demonstrations, were executed on the various robots. In line with the spirit of LfO, an effort was made to create a system that is as independent on the robot’s hardware as possible. Attention was directed towards hand motions to design skill agents that move the hand to satisfy conditions derived from the task requirements and physical constraints from the environment. By taking hand motions as a reference, differences in hardware among robots are absorbed into inverse kinematics (IK), allowing the easy substitution of an appropriate IK when hardware changes, ensuring many skill agents can be used.
In environments with physical constraints, adjustments to hand trajectories are often necessary; such a concept is known as compliant manipulation (Mason (1981)). Recently, reinforcement learning (RL) has been applied to address household tasks, including compliant manipulation. However, traditional RL methods have primarily focused on designing policies for specific tasks and requiring individual learning for each new task. We proposed policies that consider constraints applicable to various tasks by grouping multiple tasks based on the types of physical constraints. The type of physical constraints are determined by the characteristics of the imposed force directions. Consequently, a general policy is learned based on these characteristics.
End-to-end generation of robot programs using large language models (LLM) has been proposed in general, such as Vemprala et al. (2023); Wake et al. (2023); Yu et al. (2023). However, many robot tasks require compliant manipulation. In the case of compliant manipulation, adjusting the hand trajectory using force feedback becomes necessary. While systems based on LLMs can provide an overview of program design, generating programs with adjusting capabilities using local feedback is challenging. To address this issue, we proposed preparing a standard set of machine-independent executable skill agents using reinforcement learning or similar methods. This standard set can be provided as a prompt to the LLMs, enabling the design of a robot system that can actually execute. We consider this approach holds promise for resolving the challenge of incorporating local feedback in robot program generation using LLMs.
In this paper, we utilized Shadow Hand on two different robots. As a result, we were able to share the same grasp skill-agent library. In the case of the parallel gripper of Fecth, we resolved the difference in the hardware by only changing the grasp skill-agent library. Standardizing grasp skill agents across different robotic hands remains a challenge for future research.